How much coffee worth $12 per pound should be mixed with coffee worth $8 per pound to produce 50 pounds of a coffee blend worth $8.80 per pound?
Answer by checkley71(8403)

12X+8(50-X)=8.8*50
12X+400-8X=440
4X=440-400
4X=40
X=40/4
X=10 POUNDS OF $12 COFFEE.
50-10=40 POUNDS OF $8 COFFEE.
PROOF
12*10+8*40=8.8*50
120+320=440
440=440